# New results for the Cahn-Hilliard equation with non-degenerate mobility: well-posedness and longtime behavior

### v1

@@ 2.1. Notation and function spaces
$\begin{split}&\frac{1}{2}\frac{{\rm d}}{{\rm d}t}\int_{\Omega}b(\varphi)\left|\nabla\mu\right|^{2}\,{\rm d}x+\frac{1}{2}\left\|\nabla\partial_{t}\varphi\right\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}F^{\prime\prime}(\varphi)\left|\partial_{t}\varphi\right|^{2}\,{\rm d}x\\ &\leq C\left\|\partial_{t}\varphi\right\|_{L^{2}(\Omega)}^{2}+\frac{C}{\tau}\left(\|\nabla\mu\|_{L^{2}(\Omega)}^{4}+\|\nabla\mu\|_{L^{2}(\Omega)}^{4+\frac{4\tau}{1+\tau}}+\|\nabla\mu\|_{L^{2}(\Omega)}^{4+\frac{4\tau}{1-\tau}}\right)\!,\end{split}$ (4.13)
@@
$\begin{split}B(0)&\leq C\left(1+\left\|\nabla\mu(0)\right\|_{L^{2}(\Omega)}^{2}\right)\leq C\left(1+\left\|\Delta\varphi_{0}+F^{\prime}(\varphi_{0})\right\|_{H^{1}(\Omega)}^{2}\right)<\infty.\end{split}$ (4.18)
@@
It is apparent that (4.20) only provides a local bound in time on $B$, where $T_{\star}$ depends on the norm of the initial data. To overcome this issue, we propose an alternative argument leading to a global bound in time on $B$. We report the following Gronwall type result proved in [36, Lemma 2.1]:
@@
$\sup_{0\leq t\leq T}\int_{\Omega}|\nabla\mu(t)|^{2}\,{\rm d}x\leq\frac{2}{b_{m}}\left(C\left(1+\left\|\Delta\varphi_{0}+F^{\prime}(\varphi_{0})\right\|_{H^{1}(\Omega)}^{2}\right)\right)^{{\rm exp}\left(C(1+T)\right)}\!,\quad\forall\,T>0.$ (4.25)
@@
$\sup_{0\leq t\leq 1}\int_{\Omega}|\nabla\mu(t)|^{2}\,{\rm d}x\leq\frac{C}{b_{m}}\left(1+\left\|\Delta\varphi_{0}+F^{\prime}(\varphi_{0})\right\|_{H^{1}(\Omega)}^{2}\right)\!.$ (4.28)
In order to derive a global control independent of time, we report the uniform Gronwall lemma (see [44, Chapter III, Lemma 1.1])
@@
We estimate the right hand side as follows. By standard computations, we find
@@
In light of (3.13) and (4.32), the latter implies that $\mu\in L^{2}_{\operatorname{uloc}}([0,\infty);H^{3}(\Omega))$. Since the uniqueness of the strong solutions follows from [5, Theorem 2.2] (cf. Theorem 1.1 - part (2) and [41, Theorem 2.3]), the proof of Proposition 4.1 is complete. ∎
@@
Let $\varphi$ be the global weak solutions originating from the initial condition $\varphi_{0}$. Fix $\tau>0$. Let $I=\{t\in(0,1]:\partial_{\textbf{{n}}}\phi(t)=0\text{ on }\partial\Omega\}$. We recall that $I$ has full Lebesgue measure . Since
@@
$V_{m}(\Omega)=\Big{\{}f\in H_{(m)}^{1}(\Omega)\cap L^{\infty}(\Omega):\|f\|_{L^{\infty}(\Omega)}\leq 1\Big{\}},$
@@ 4. $\diamond$
$\exists\,E_{\infty}:\lim_{t\to\infty}E(\varphi(t))\quad\text{and}\quad E(\varphi^{\prime})=E_{\infty},\ \forall\,\varphi^{\prime}\in\omega(\varphi_{0}),$ (5.1)
@@
$|\widetilde{E}(\varphi)-\widetilde{E}(\varphi^{\prime})|^{1-\theta}\leq C\left\|D\widetilde{E}(\varphi)\right\|_{H^{1}_{(0)}(\Omega)^{\prime}}\!,\quad\forall\,\varphi\in H^{1}_{(m)}(\Omega):\,\|\varphi-\varphi^{\prime}\|_{H^{1}_{(0)}}\leq\beta.$ (5.4)
Since $\omega(\varphi_{0})$ is compact in $H^{1}_{(m)}(\Omega)$, we cover $\omega(\varphi_{0})$ by finitely many open balls $\{B_{i}\}_{i=1}^{N}$ in $H^{1}_{(m)}(\Omega)$ centered in $\varphi^{\prime}_{i}\in\omega(\varphi_{0})$ with radius $\beta_{i}$, where $\beta_{i}$ is the constant from Lemma 5.1 corresponding to $\varphi^{\prime}_{i}$. Recalling that $\widetilde{E}|_{\omega(\varphi_{0})}=E_{\infty}$ and set $U:=\bigcup_{i=1}^{N}B_{i}$, there exist universal constants $\widetilde{\theta}\in\left(0,\frac{1}{2}\right)$ and $\widetilde{C}>0$ such that
@@
$\left\|\mathcal{G}_{q}f\right\|_{H^{2}(\Omega)}\leq\frac{sC}{2}\left\|f\right\|_{L^{2}(\Omega)}+C^{\frac{s}{2}}\left(\frac{2s^{2}}{s-2}\right)^{\frac{s}{4}}\left\|\nabla q\right\|_{L^{2}(\Omega)}^{\frac{s-2}{2}}\left\|q\right\|_{H^{2}(\Omega)}\left\|\nabla\mathcal{G}_{q}f\right\|_{L^{2}(\Omega)}.$ (A.1)
Acknowledgments. AG is supported by the MUR grant Dipartimento di Eccellenza 2023-2027 of Dipartimento di Matematica, Politecnico di Milano. The four authors are supported by Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), Istituto Nazionale di Alta Matematica (INdAM).

### v2

@@ 2.1. Notation and function spaces
$\begin{split}&\frac{1}{2}\frac{{\rm d}}{{\rm d}t}\int_{\Omega}b(\varphi)\left|\nabla\mu\right|^{2}\,{\rm d}x+\frac{1}{2}\left\|\nabla\partial_{t}\varphi\right\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}F^{\prime\prime}(\varphi)\left|\partial_{t}\varphi\right|^{2}\,{\rm d}x\\ &\quad\leq C\left\|\partial_{t}\varphi\right\|_{L^{2}(\Omega)}^{2}+\frac{C}{\tau}\left(\|\nabla\mu\|_{L^{2}(\Omega)}^{4}+\|\nabla\mu\|_{L^{2}(\Omega)}^{4+\frac{4\tau}{1+\tau}}+\|\nabla\mu\|_{L^{2}(\Omega)}^{4+\frac{4\tau}{1-\tau}}\right)\!,\end{split}$ (4.13)
@@
$\begin{split}B(0)&\leq C\left(1+\left\|\nabla\mu(0)\right\|_{L^{2}(\Omega)}^{2}\right)\leq C\left(1+\left\|-\Delta\varphi_{0}+F^{\prime}(\varphi_{0})\right\|_{H^{1}(\Omega)}^{2}\right)<\infty.\end{split}$ (4.18)
@@
It is apparent that (4.20) only provides a local bound in time on $B$, where $T_{\star}$ depends on the norm of the initial data. To overcome this issue, we propose an alternative argument leading to a global bound in time on $B$. We report the following Gronwall type result proved in [35, Lemma 2.1]:
@@
$\sup_{0\leq t\leq T}\int_{\Omega}|\nabla\mu(t)|^{2}\,{\rm d}x\leq\frac{2}{b_{m}}\left(C\left(1+\left\|-\Delta\varphi_{0}+F^{\prime}(\varphi_{0})\right\|_{H^{1}(\Omega)}^{2}\right)\right)^{{\rm exp}\left(C(1+T)\right)}\!,\quad\forall\,T>0.$ (4.25)
@@
$\sup_{0\leq t\leq 1}\int_{\Omega}|\nabla\mu(t)|^{2}\,{\rm d}x\leq\frac{C}{b_{m}}\left(1+\left\|-\Delta\varphi_{0}+F^{\prime}(\varphi_{0})\right\|_{H^{1}(\Omega)}^{2}\right)\!.$ (4.28)
In order to derive a global control independent of time, we report the uniform Gronwall lemma (see [43, Chapter III, Lemma 1.1])
@@
We estimate the right-hand side as follows. By standard computations, we find
@@
In light of (3.13) and (4.32), the latter implies that $\mu\in L^{2}_{\operatorname{uloc}}([0,\infty);H^{3}(\Omega))$. Since the uniqueness of the strong solutions follows from [5, Theorem 2.2] (cf. Theorem 1.1 - part (2) and [40, Theorem 2.3]), the proof of Proposition 4.1 is complete. ∎
@@
Let $\varphi$ be the global weak solutions originating from the initial condition $\varphi_{0}$. Fix $\tau>0$. Let $I=\{t\in(0,1]:\partial_{\textbf{{n}}}\varphi(t)=0\text{ on }\partial\Omega\}$. We recall that $I$ has full Lebesgue measure . Since
@@
$V_{m}=\Big{\{}f\in H_{(m)}^{1}(\Omega)\cap L^{\infty}(\Omega):\|f\|_{L^{\infty}(\Omega)}\leq 1\Big{\}},$
@@ 4. $\diamond$
$\exists\,E_{\infty}:=\lim_{t\to\infty}E(\varphi(t))\quad\text{and}\quad E(\varphi^{\prime})=E_{\infty},\ \forall\,\varphi^{\prime}\in\omega(\varphi_{0}),$ (5.1)
@@
$|\widetilde{E}(\varphi)-\widetilde{E}(\varphi^{\prime})|^{1-\theta}\leq C\left\|D\widetilde{E}(\varphi)\right\|_{H^{1}_{(0)}(\Omega)^{\prime}}\!,$ (5.4)
for all $\varphi\in H^{1}_{(m)}(\Omega)$ such that $\|\varphi-\varphi^{\prime}\|_{H^{1}_{(0)}(\Omega)}\leq\beta$.
Since $\omega(\varphi_{0})$ is compact in $H^{1}_{(m)}(\Omega)$, we cover $\omega(\varphi_{0})$ by finitely many open balls $\{B_{i}\}_{i=1}^{N}$ in $H^{1}_{(m)}(\Omega)$ centered at $\varphi^{\prime}_{i}\in\omega(\varphi_{0})$ with radius $\beta_{i}$, where $\beta_{i}$ is the constant from Lemma 5.1 corresponding to $\varphi^{\prime}_{i}$. Recalling that $\widetilde{E}|_{\omega(\varphi_{0})}=E_{\infty}$ and set $U:=\bigcup_{i=1}^{N}B_{i}$, there exist universal constants $\widetilde{\theta}\in\left(0,\frac{1}{2}\right]$ and $\widetilde{C}>0$ such that
@@
$\left\|\mathcal{G}_{q}f\right\|_{H^{2}(\Omega)}\leq\frac{sC}{2}\left\|f\right\|_{L^{2}(\Omega)}+C^{\frac{s}{2}}\left(\frac{2s^{2}}{s-2}\right)^{\frac{s}{4}}\left\|\nabla q\right\|_{L^{2}(\Omega)}^{\frac{s-2}{2}}\left\|q\right\|_{H^{2}(\Omega)}\left\|\nabla\mathcal{G}_{q}f\right\|_{L^{2}(\Omega)}.$
Acknowledgments. The authors are grateful to the referee for the helpful comments that improved the presentation of this paper. A. Giorgini is supported by the MUR grant Dipartimento di Eccellenza 2023-2027 of Dipartimento di Matematica, Politecnico di Milano. This work is supported by Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), Istituto Nazionale di Alta Matematica (INdAM).
Competing Interests and funding. The authors do not have any financial or non-financial interests that are directly or indirectly related to the work submitted for publication.
Data availability statement. No further data is used in this manuscript.